# ReMoD: Rethinking Modality Contribution in Multimodal Stance Detection via Dual Reasoning

### v1

@@ Introduction
Deliberate Reflective Reasoning is analogous to the ‘System 2’ slow thinking process in dual-process cognitive theory and serves as the refinement module of our framework. This process is governed by experience formulation, where novel insights are generated through structured reflection, and experience evolution, where the experience pools are refined by integrating these new insights.
Experiences Formulation.
@@
The primary objective of experience formulation is to facilitate metacognitive reflection, enabling the agent to distill actionable experience from its performance outcomes. To this end, the ReMoD agent employs a structured, multi-step reasoning process, analogous to a CoT, to generate new insights.
First, to generate the modality experience, ReMoD evaluates the accuracy of its initial stance predictions across text-only $\hat{y}_{u}$, vision-only $\hat{y}_{\mathcal{V}}$, and combined multimodal $\hat{y}_{(u,\mathcal{V})}$ contexts, along with their corresponding rationales $r_{u}$, $r_{\mathcal{V}}$, and $r_{(u,\mathcal{V})}$. The outcomes are then distilled into a high-level modality insight $I_{m}$. This insight functions as a strategic heuristic, guiding the optimal balance of unimodal and multimodal cues to improve stance detection performance. In parallel, forming semantic experience involves ReMoD initiating deep reflection on its reasoning process, with guidance from the initial prediction outcomes of the same text-only, vision-only, multimodal contexts, and the corresponding rationales. This reflective process involves a joint analysis of the source text knowledge $\mathcal{K}_{u}$ and visual knowledge $\mathcal{K}_{v}$ to scrutinize the depth and comprehensiveness of the reasoning that led to those predictions. These reflections are subsequently distilled into a semantic insight $I_{s}$, which articulates actionable strategies for deepening semantic understanding for the task.
Experiences Evolution.
@@
Upon generating new insights $I_{m}$ and $I_{s}$, ReMoD updates its experience pools to consolidate and refine existing experiences by first employing the modality-synergistic retrieval mechanism to retrieve the top-$k$ most relevant prior experiences from each pool. These experiences, denoted as $\{\mathcal{\bar{S}}_{i}^{\text{ME}}\}^{k}_{i=1}$ from MEP and $\{\mathcal{\bar{S}}_{i}^{\text{SE}}\}^{k}_{i=1}$ from SEP, have scores exceeding a relevance threshold $\tau$ and are selected based on the current text information $u$ and visual information $\mathcal{V}$. If no experiences meet this threshold, the new insights are encapsulated as new key-value entries and appended to their respective pools. Conversely, if relevant experiences are found, ReMoD performs targeted updates on these entries to prevent redundant memories.
@@
This evolutionary update hinges on a fusion operation. For each retrieved modality experience, ReMoD initiates a re-reasoning process, using a prompt to synthesize the prior experience $\{\mathcal{\bar{S}}^{\text{ME}}_{i}\}^{k}_{i=1}$ with the new modality insight $I_{m}$. This generates a more comprehensive, updated modality experience pool $\mathcal{\hat{S}}^{\text{ME}}$, which then overwrites the original entry. A parallel process is executed for the semantic experiences, updating $\{\mathcal{\bar{S}}_{i}^{\text{SE}}\}^{k}_{i=1}$ with $I_{s}$ to produce new semantic experience pool $\mathcal{\hat{S}}_{i}^{\text{SE}}$. This dual-update ensures both experience pools remain synchronized and continuously refined. In essence, this mechanism enables ReMoD to build a compact and coherent knowledge base through continuous refinement, supporting efficient and scalable lifelong learning.
@@
Table 1: In-target multimodal stance detection results (%). Best and second-best scores are in bold and underline, respectively. indicates our ReMod model significantly outperforms baselines ($p<0.05$). The dashed line separates fine-tuned from non-fine-tuned methods. w/ denotes the backbone, and $\uparrow$ indicates improvement over the original MLLM.
Stance Inference.
@@
During the inference phase, ReMoD first leverages the SAM to extract visual information $\mathcal{V}$ from the input image, capturing fine-grained visual manifestations of semantic entities. This visual information is then paired with the text input $u$ to form a multimodal context. It then employs its modality-synergistic retrieval mechanism to fetch the relevant prior experiences from both the modality and semantic experience pools. Finally, ReMoD integrates these experiences with multimodal context and the target to derive the final stance prediction, ensuring decisions are based on immediate data and guided by its refined long-term understanding.
Experiments
Experiment Settings
Dataset and Metrics.
@@
To rigorously evaluate our ReMoD, we benchmark its performance on the MMSD dataset (1-liang2024multi), which comprises five distinct domains: Multi-modal COVID-CQ (MCCQ), Multi-modal Russo-Ukrainian Conflict (MRUC), Multi-modal Taiwan Question (MTWQ), Multi-modal Twitter Stance Election 2020 (MTSE), and Multi-modal Will-They-Won’t-They (MWTWT). Following the MMSD dataset, we adopt the Macro-F1 score as the primary metric for evaluation.
Implementation Details.
@@
Our ReMoD framework is implemented upon the Qwen-2.5-VL-32B model as its backbone. For the experience pool retrieval, we set the coefficient $\alpha$ to 0.7, a relevance threshold $\tau$ to 0.8, and retrieve the top kk=3 experiences to balance context with noise. All input images are resized to a minimum of 28 pixels on their shorter side to meet requirements, while entity-related knowledge is sourced from Wikipedia using the Python Wikipedia library. More details are shown in the appendix.
Baseline Methods.
@@
We evaluate our method against three categories of baselines: Text-only baselines: 1) BERT (33-devlin2019bert), using uncased BERT-base model; 2) RoBERTa (34-liu2019roberta), using RoBERTa-base model; 3) KEBERT (35-kawintiranon2022polibertweet); 4) LLaMA2 (36-touvron2023llama), using LLaMA2-70B-chat; 5) GPT4. Vision-only baselines: 1) ResNet (37-he2016deep), using ResNet-50 v1.5; 2) ViT (38-dosovitskiy2020image), using ViT-base-patch16-224 model; 3) Swin Transformer (SwinT) (39-liu2021swin), using SwinV2-base-patch4-window12-192-22k model. Multimodal baselines: 1) ViLT (40-kim2021vilt), using vilt-b32-mlm model; 2) CLIP (41-radford2021learning), using CLIP-ViT-base-patch32 model; 3) BERT+ViT, a model that uses BERT and ViT as the text and image encoders; 4) Qwen-VL (42-bai2023qwen), using the Qwen-VL-Chat7B. 5) Qwen2.5-VL (43-bai2025qwen2), using the Qwen2.5-VL-32B-Instruct; 6) GPT4-Vision ; 7) MiMo (44-xiaomi2025mimo), using MiMo-7B-RL, 8) LLaVa (45-li2024llava), using LLaVa-v1.6-Mistral-7B; 9) TMPT and TMPT+CoT(1-liang2024multi).
@@
Table 2: Zero-shot multimodal stance detection results (%). Best and second-best scores are in bold and underline, respectively. indicates our ReMod model significantly outperforms baselines ($p<0.05$). The dashed line separates fine-tuned from non-fine-tuned methods. w/ denotes the backbone, and $\uparrow$ indicates improvement over the original MLLM.
Main Results
In-target Multimodal Stance Detection.
As presented in Table 1, our ReMoD achieves competitive performance across most baselines, even powerful LLMs such as GPT4 and GPT4-Vision in multiple scenarios. While TMPT+CoT shows stronger results on some targets, it relies on training with GPT4-generated CoT examples. Instead, ReMoD highlights its training-free superiority by leveraging accumulated experience for adaptive modality contribution rethinking.
A deeper analysis of baseline performance reveals the intricate challenges in multimodal fusion. The consistent outperformance of text-only models over vision-only counterparts confirms that text serves as the primary modality for stance determination. Yet, the superior performance of advanced multimodal models like GPT4-Vision compared to their text-only variants validates the supplementary value of visual cues, underscoring the necessity of cross-modality integration. Notably, the inconsistent performance gains across fusion approaches, combined with the struggles of simpler frameworks like CLIP and TMPT, underscore a critical deficiency, as these models fail to handle the varying contribution of each modality on a per-sample basis, allowing noisy or irrelevant visual data to degrade overall performance. In contrast, ReMoD leverages its dual-reasoning paradigm to dynamically rethink modality contributions, thereby enhancing its final predictive accuracy.
Zero-shot Multimodal Stance Detection.
As shown in Table 2, the results expose the inherent fragility of models reliant on in-domain training. While smaller architectures like BERT and RoBERTa suffer a precipitous performance collapse, even powerful LLMs such as GPT4 and GPT4-Vision demonstrate significant performance volatility when confronted with unseen targets. This underscores a fundamental limitation in current paradigms: true zero-shot generalization remains a formidable and largely unsolved challenge. In this demanding scenario, ReMoD establishes a new standard for robustness and adaptability. It achieves state-of-the-art results on a majority of the evaluated datasets. This stability, particularly when contrasted with the erratic outputs of leading LLM-based baselines, provides compelling evidence of a superior generalization capability.
@@
Ablation Study
To evaluate the impact of each component within our ReMoD, we conduct an ablation study, with results reported in Table Zero-shot Multimodal Stance Detection.. The findings unequivocally demonstrate the synergistic importance of all modules, as the full ReMoD model consistently outperforms any ablated version across both in-target and zero-shot settings. The most critical components are the experience pools. Without using modality experience (w/o MEP) results in the most significant performance degradation, particularly in the zero-shot scenario, highlighting the indispensable role of assessing cross-sample modality contribution for robust generalization. Similarly, the exclusion of using semantic experience (w/o SEP) causes a severe decline, confirming the necessity of leveraging semantic context for accurate stance inference. Finally, while removing the CoT (w/o CoT) module also impairs performance, the impact is less pronounced, suggesting that CoT functions as a higher-level reasoning synthesizer, which organizes and refines the foundational evidence provided by the MEP and SEP to formulate a coherent final stance.
@@
Effect of Different Backbones
We assess the performance of ReMoD with different MLLM backbones in both in-target and zero-shot settings, as detailed in Table 1 and Table 2. Smaller backbones generally yield weaker performance, likely due to their limited visual perception capabilities. However, integrating any MLLM as ReMoD’s backbone still brings significant improvements marked by $\uparrow$ compared to standalone operation, with Qwen-VL showing particularly notable gains. This consistent enhancement across diverse architectures highlights that our framework effectively handles varying modality contributions by dynamically weighting them based on their actual expressive power for target stances. It thus compensates for limitations in smaller backbones, ensuring robust performance regardless of the backbone’s inherent biases toward text or vision.
Different Relevance Thresholds in Retrieval
@@
We analyzed the effect of varying relevance thresholds (0.1, 0.3, 0.5, 0,7, 0.8, and 0.9) on retrieval performance for both in-target and zero-shot settings, as shown in Figure 3. The results indicate that in both settings, performance tends to be weaker at lower thresholds and gradually degrades once the threshold exceeds 0.8. This trend highlights a critical trade-off between the quality and quantity of retrieved experiences. A low threshold risks retrieving irrelevant or poor-quality experiences, which can introduce noise and impair the reasoning process. Conversely, an excessively high threshold becomes too restrictive, retrieving an insufficient number of experiences and thus limiting the model’s ability to leverage its accumulated knowledge for robust inference.
@@
Figure 4: Case study between our ReMoD and directly CoT method based on the same LLM backbone.
Case Study
@@
To further underscore the strengths of ReMoD’s dual reasoning, we analyze two examples in Figure 4 by comparing the results of direct CoT with ReMoD. In Example (a), direct CoT literally interprets ‘stan Putin’ in the text as support but ignores the negative context of ‘cringe’, which renders ‘stan’ sarcastic. ReMoD’s modality experience emphasizes weighting textual evidence via contextual judgment, enabling recognition of implicit sarcasm through the contradiction between cringe and stan to avoid being misled by isolated explicit terms. Its semantic experience clarifies that such sarcasm criticizes Russia-related phenomena, thus linking the text to the target stance of opposing Russia.
@@
For Example (b), direct CoT infers support for the drug from surface information like India lifting the export ban in text and leaders’ collaboration in images. In contrast, ReMoD’s modality experience stresses prioritizing explicit stance information across modalities, equal weighting, and default neutrality, while it deepens understanding via common sense and geopolitical context, revealing that the text describes objective policy adjustments amid international games rather than the author’s subjective endorsement of the drug’s efficacy. By curbing surface information misinterpretation and calibrating reasoning with background knowledge, ReMoD accurately captures the neutral stance.
Conclusion
@@
This paper proposes ReMoD, a framework for multimodal stance detection that employs a dual-reasoning paradigm inspired by human cognition. By integrating experience-driven intuitive reasoning with deliberate reflective reasoning, it dynamically captures the variable contributions of different modalities, overcoming the limitations of fixed fusion strategies. The dual-pool mechanism enables the accumulation and refinement of modality and semantic insights, enhancing robustness and context awareness. Experiments on MMSD show ReMoD outperforms most baselines in both in-target and zero-shot settings, with strong generalization.

### v2

@@ 3.3 Meta-cognitive Reflective Reasoning
Meta-cognitive Reflective Reasoning is analogous to the ‘System 2’ slow thinking process in dual-process cognitive theory and serves as the refinement module of our framework. This process is governed by experience formulation, where novel insights are generated through structured reflection, and experience evolution, where the experience pools are refined by integrating these new insights.
3.3.1 Experiences Formulation
@@
The primary objective of experience formulation is to facilitate meta-cognitive reflection, enabling the agent to distill actionable experience from its performance outcomes. To this end, the MIND agent employs a structured, multi-step reasoning process, analogous to a CoT, to generate new insights.
To generate the modality experience, the Modality-CoT initiates a self-diagnostic process. It evaluates the accuracy of the initial predictions for text $\hat{y}_{u}$, vision $\hat{y}_{\mathcal{V}}$, and combined multimodal input $\hat{y}_{(u,\mathcal{V})}$, along with their corresponding rationales $r_{u}$, $r_{\mathcal{V}}$, and $r_{(u,\mathcal{V})}$. By analyzing the failures and successes of these initial judgments, it distills a high-level modality insight $\mathcal{I}_{m}$. This insight is not merely a summary, but an adaptive fusion strategy that dictates how to dynamically balance unimodal and multimodal cues in future reasoning.
In parallel, the Semantic-CoT is tasked with forming the semantic experience. This process involves a deep interrogation of the initial reasoning’s semantic grounding. It synthesizes the agent’s internal rationales with the externally-grounded textual knowledge $\mathcal{K}_{u}$ and visual knowledge $\mathcal{K}_{v}$. This synthesis allows the agent to identify and rectify content-level biases or shallow interpretations. The outcome is a semantic insight $\mathcal{I}_{s}$, an actionable directive that refines the model’s strategies for achieving deeper semantic and contextual understanding.
3.3.2 Experiences Evolution
@@
Upon deriving new insights $\mathcal{I}_{m}$ and $\mathcal{I}_{s}$, MIND updates its experience pools to consolidate and refine existing experiences by first employing the modal synergistic retrieval mechanism to retrieve the top-$k$ most relevant prior experiences from each pool. These experiences, denoted as $\{\mathcal{\bar{S}}_{i}^{\text{ME}}\}^{k}_{i=1}$ from MEP and $\{\mathcal{\bar{S}}_{i}^{\text{SE}}\}^{k}_{i=1}$ from SEP, have scores exceeding a relevance threshold $\tau$ and are selected based on the current text information $u$ and visual information $\mathcal{V}$. If no experiences meet this threshold, the new insights are encapsulated as new key-value entries and appended to their respective pools. Conversely, MIND performs targeted updates on these entries to prevent redundant memories.
@@
This evolutionary update hinges on a fusion operation. For each retrieved modality experience, MIND initiates a re-reasoning process, using a prompt to synthesize the prior experience $\{\mathcal{\bar{S}}^{\text{ME}}_{i}\}^{k}_{i=1}$ with the new modality insight $I_{m}$. This generates a more comprehensive, updated modality experience pool $\mathcal{\hat{S}}^{\text{ME}}$, which then overwrites the original entry. A parallel process is executed for the semantic experiences, updating $\{\mathcal{\bar{S}}_{i}^{\text{SE}}\}^{k}_{i=1}$ with $I_{s}$ to produce new semantic experience pool $\mathcal{\hat{S}}_{i}^{\text{SE}}$. This dual-update ensures both experience pools remain synchronized and continuously refined. In essence, this mechanism enables MIND to build a compact and coherent knowledge base through continuous refinement, supporting efficient learning.
@@
Table 1: In-target results (%). Best and second-best scores are in bold and underline, respectively. indicates our MIND model significantly outperforms baselines ($p<0.05$). The dashed line separates fine-tuned from non-fine-tuned methods. w/ denotes the backbone, and $\uparrow$ indicates improvement over the original MLLM.
3.3.3 Stance Inference
@@
During the inference phase, MIND first leverages the SAM to extract visual information $\mathcal{V}$ from the input image, capturing fine-grained visual manifestations of semantic entities. This visual information is then paired with the text input $u$ to form a multimodal context. It then employs its modality-synergistic retrieval mechanism to fetch the relevant prior experiences from both the modality and semantic experience pools. Finally, MIND integrates these experiences with multimodal context and the target to derive the final stance prediction, ensuring decisions are based on immediate data and guided by its refined long-term understanding.
4 Experiments
4.1 Experiment Settings
@@
Dataset and Metrics. To evaluate our MIND, we benchmark its performance on the MMSD dataset Liang et al. (2024), which comprises five distinct domains: Multi-modal COVID-CQ (MCCQ), Multi-modal Russo-Ukrainian Conflict (MRUC), Multi-modal Taiwan Question (MTWQ), Multi-modal Twitter Stance Election 2020 (MTSE), and Multi-modal Will-They-Won’t-They (MWTWT). Following the MMSD dataset, we adopt the Macro-F1 score as the primary metric for evaluation.
@@
Implementation Details. Our MIND method is implemented upon the Qwen-2.5-VL-32B model as its backbone. For the experience pool retrieval, we set the coefficient $\alpha$ to 0.7, a relevance threshold $\tau$ to 0.7, and retrieve the top $k=3$ experiences to balance context with noise. Entity-related knowledge is sourced from Wikipedia using the Python Wikipedia library.
4.2 Baseline Methods.
@@
We evaluate our method against three categories of baselines: Text-only baselines: 1) BERT Devlin et al. (2019), using uncased BERT-base model; 2) RoBERTa Liu et al. (2019), using RoBERTa-base model; 3) KEBERT Kawintiranon and Singh (2022); 4) LLaMA2 Touvron et al. (2023), using LLaMA2-70B-chat; 5) GPT4. Vision-only baselines: 1) ResNet He et al. (2016), using ResNet-50 v1.5; 2) ViT Dosovitskiy et al. (2020), using ViT-base-patch16-224 model; 3) Swin Transformer (SwinT) Liu et al. (2021), using SwinV2-base-patch4-window12-192-22k model. Multimodal baselines: 1) ViLT Kim et al. (2021), using vilt-b32-mlm model; 2) CLIP Radford et al. (2021), using CLIP-ViT-base-patch32 model; 3) BERT+ViT, a model that uses BERT and ViT as the text and image encoders; 4) Qwen-VL Bai et al. (2023), using the Qwen-VL-Chat7B. 5) Qwen2.5-VL Bai et al. (2025), using the Qwen2.5-VL-32B-Instruct; 6) GPT4-Vision ; 7) MiMo Xiaomi et al. (2025), using MiMo-7B-RL, 8) LLaVa Li et al. (2024), using LLaVa-v1.6-Mistral-7B; 9) TMPT and TMPT+CoT Liang et al. (2024).
@@
Table 2: Zero-shot results (%). Best and second-best scores are in bold and underline, respectively. indicates our MIND model significantly outperforms baselines ($p<0.05$). The dashed line separates fine-tuned from non-fine-tuned methods. w/ denotes the backbone, and $\uparrow$ indicates improvement over the original MLLM.
4.3 Main Results
4.3.1 In-target Multimodal Stance Detection.
As presented in Table 1, our MIND achieves competitive performance across most baselines, even powerful LLMs such as GPT4 and GPT4-Vision in multiple scenarios. While TMPT+CoT shows stronger results on some targets, it relies on training with GPT4-generated CoT examples. Instead, MIND highlights its superiority by leveraging accumulated experience for adaptive modality contribution rethinking.
A deeper analysis of baseline performance reveals the intricate challenges in multimodal fusion. The consistent outperformance of text-only models over vision-only counterparts confirms that text serves as the primary modality for stance determination. Yet, the superior performance of advanced multimodal models like GPT4-Vision compared to their text-only variants validates the supplementary value of visual cues, underscoring the necessity of cross-modality integration. Notably, the inconsistent performance gains across fusion approaches, combined with the struggles of simpler frameworks like CLIP and TMPT, underscore a critical deficiency, as these models fail to handle the varying contribution of each modality on a per-sample basis, allowing noisy or irrelevant visual data to degrade overall performance. In contrast, MIND leverages its dual-reasoning paradigm to dynamically rethink modality contributions, thereby enhancing its final predictive accuracy.
4.3.2 Zero-shot Multimodal Stance Detection.
As shown in Table 2, the results expose the inherent fragility of models reliant on in-domain training. While smaller architectures like BERT and RoBERTa suffer a precipitous performance collapse, even powerful LLMs such as GPT4 and GPT4-Vision demonstrate significant performance volatility when confronted with unseen targets. This underscores a fundamental limitation in current paradigms: true zero-shot generalization remains a formidable and largely unsolved challenge. In this demanding scenario, MIND establishes a new standard for robustness and adaptability. It achieves state-of-the-art results on a majority of the evaluated datasets. This stability, particularly when contrasted with the erratic outputs of leading LLM-based baselines, provides compelling evidence of a superior generalization capability.
@@ 4.3 Main Results
4.4 Ablation Study
To evaluate the impact of each component within our MIND, we conduct an ablation study, with results reported in Table 4.3.2. The findings unequivocally demonstrate the synergistic importance of all modules, as the full MIND model consistently outperforms any ablated version across both in-target and zero-shot settings. The most critical components are the experience pools. Without using modality experience (w/o MEP) results in the most significant performance degradation, particularly in the zero-shot scenario, highlighting the indispensable role of assessing cross-sample modality contribution for robust generalization. Similarly, the exclusion of using semantic experience (w/o SEP) causes a severe decline, confirming the necessity of leveraging semantic context for accurate stance inference. Finally, while removing the CoT (w/o CoT) module also impairs performance, the impact is less pronounced, suggesting that CoT functions as a higher-level reasoning synthesizer, which organizes and refines the foundational evidence provided by the MEP and SEP to formulate a coherent final stance.
4.5 Effect of Different Backbones
We assess the performance of MIND with different MLLM backbones in both in-target and zero-shot settings, as detailed in Table 1 and Table 2. Smaller backbones generally yield weaker performance, likely due to their limited visual perception capabilities. However, integrating any MLLM as MIND’s backbone still brings significant improvements marked by $\uparrow$ compared to standalone operation, with Qwen-VL showing particularly notable gains. This consistent enhancement across diverse architectures highlights that our framework effectively handles varying modality contributions by dynamically weighting them based on their actual expressive power for target stances. It thus compensates for limitations in smaller backbones, ensuring robust performance regardless of the backbone’s inherent biases toward text or vision.
@@ 4.5 Effect of Different Backbones
4.6 Different Relevance Thresholds
@@
We analyzed the effect of varying relevance thresholds (0.1, 0.3, 0.5, 0,7, and 0.9) on retrieval performance for both in-target and zero-shot settings, as shown in Figure 3. The results indicate that in both settings, performance tends to be weaker at lower thresholds and gradually degrades once the threshold exceeds 0.7. This trend highlights a critical trade-off between the quality and quantity of retrieved experiences. A low threshold risks retrieving irrelevant or poor-quality experiences, which can introduce noise and impair the reasoning process. Conversely, an excessively high threshold becomes too restrictive, retrieving an insufficient number of experiences and thus limiting the model’s ability to leverage its accumulated knowledge for robust inference.
@@
Figure 4: Case study between our MIND and directly CoT method based on the same LLM backbone.
4.7 Case Study
@@
To further underscore the strengths of MIND’s dual reasoning, we analyze two examples in Figure 4 by comparing the results of direct CoT with MIND. In Example (a), direct CoT literally interprets ‘stan Putin’ in the text as support but ignores the negative context of ‘cringe’, which renders ‘stan’ sarcastic. MIND’s modality experience emphasizes weighting textual evidence via contextual judgment, enabling recognition of implicit sarcasm through the contradiction between cringe and stan to avoid being misled by isolated explicit terms. Its semantic experience clarifies that such sarcasm criticizes Russia-related phenomena, thus linking the text to the target stance of opposing Russia.
@@
For Example (b), direct CoT infers support for the drug from surface information like India lifting the export ban in text and leaders’ collaboration in images. In contrast, MIND’s modality experience stresses prioritizing explicit stance information across modalities, equal weighting, and default neutrality, while it deepens understanding via common sense and geopolitical context, revealing that the text describes objective policy adjustments amid international games rather than the author’s subjective endorsement of the drug’s efficacy. By curbing surface information misinterpretation and calibrating reasoning with background knowledge, MIND accurately captures the neutral stance.
5 Conclusion
@@
This paper proposes MIND, a framework for multimodal stance detection that employs a dual-reasoning paradigm inspired by human cognition. By integrating experience-driven intuitive reasoning with deliberate reflective reasoning, it dynamically captures the variable contributions of different modalities, overcoming the limitations of fixed fusion strategies. The dual-pool mechanism enables the accumulation and refinement of modality and semantic insights, enhancing robustness and context awareness. Experiments on MMSD show MIND outperforms most baselines in both in-target and zero-shot settings, with strong generalization.
Limitations
A primary limitation of this work is its computational cost and inference latency. To emulate the complex dual-reasoning process of human cognition, our MIND framework involves multiple calls to the large multimodal model at various stages, including knowledge perception, intuitive reasoning, and reflective reasoning. While this deliberate design is crucial for achieving dynamic and adaptive reasoning, it inevitably leads to significant computational overhead, making it less suitable for real-time applications. Future work could explore techniques such as knowledge distillation or model compression to create a more lightweight version of the model while preserving its core reasoning capabilities.